At what time after 12:00 o’clock midnight will the minute

hand and the hour hand of a clock be on a straight line for

the first time?


1
Expert's answer
2021-09-20T16:44:43-0400

Since the Complete Angle is "360^{\circ}"; therefore, the minute hand travels "360^{\circ}"in one hour.

The clock is divided into "12" parts (hour); therefore, the hour hand travels "\frac{360^{\circ}}{12}=30^{\circ}" in one hour.

Let "x" be the position of the minute hand on the clock (in minutes)

Since one hour has "60" minutes; therefore, one hour can be expressed as "\frac{x}{60}".

So to find the position of the minute hand on the clock substitute into the relation:

Min hand in degrees - Hour hand in degrees = "180^{\circ}"

"\frac{x}{60}\cdot 360-\frac{x}{60}\cdot 30=180^{\circ}"

Simplify each fraction:

"6x-\frac{1}{2}x=180"

Subtract the terms:

"5.5x=180"

Divide both sides by "5.5" :

"x\approx 32.73"

So, at "12:32:73" both pm, the hands of a clock will be in a straight line.